OpenStudy (anonymous):

360=111+44x+(1+33x)+47x

OpenStudy (mertsj):

\[360=111+44x+1+33x+47x\] \[360=112+124x\] \[248=124x\] \[2=x\]

OpenStudy (anonymous):

Thanks!